1/F = 1/S + 1/U Solve for S

Answer by checkley75(3666)

1/F=1/S+1/U
1/S=1/F-1/U
1/S=(U-F)/FU NOW CROSS MULTIPLY
S(U-F)=FU NOW DIVIDE BOTH SIDES BY (U-F)
S=FU/(U-F)